# An analogue of Reed’s conjecture for digraphs

### v2

@@ 4 Hitting every maximum biclique with an acyclic set of vertices
Let $G$ be a graph with maximum degree $\Delta(G)\leq 3k-1$. Haxell [20] proved that, if the vertices of $G$ can be partitionned into $r$ cliques $C_{1},\dots,C_{r}$ of size $\geq 2k$, then $G$ contains an independent set $S$ of size $r$. In particular, $S$ intersects each clique $C_{i}$.
Using this result, Rabern [30] proved that every graph $G$ satisfying $\omega(G)\geq\frac{3}{4}(\Delta(G)+1)$ contains an independent set intersecting every maximum clique of $G$. This was later improved to $\omega(G)>\frac{2}{3}(\Delta(G)+1)$ by King [23], which is best possible. Christofides, Edwards, and King [10] later characterised exactly the graphs $G$ satisfying $\omega(G)=\frac{2}{3}(\Delta(G)+1)$ that do not admit an independent set intersecting every maximum cliques. Combined with King’s earlier result, they thus obtained the following. Given two graphs $G$ and $G^{\prime}$, we denote by $G\circ G^{\prime}$ the lexicographic product of $G$ by $G^{\prime}$, which is the graph with vertex-set $V(G)\times V(G^{\prime})$ and edge-set $\{(u,u^{\prime})(v,v^{\prime})\mid uv\in E(G)\textbf{~{}or~{}}u=v\text{~{}and~{}}u^{\prime}v^{\prime}\in E(G^{\prime})\}$.
Theorem 4.1 (Christofides et al. [10]).
Let $G$ be a connected graph with $\omega(G)\geq\frac{2}{3}(\Delta(G)+1)$. Then $G$ has an independent set $I$ such that $\omega(G-I)=\omega(G)-1$, unless $G=C_{n}\circ K_{p}$ with $n\geq 5$ odd and $p=\frac{1}{2}\omega(G)$.
The goal of this section is to prove the following generalisation of Theorem 4.1. Note that, in the proof of Theorem 1.6, we will only use Theorem 4.2 with ${\omega}(D)$ being larger than $\frac{2}{3}(\Delta+1)$. We find the characterisation of the equality case of independent interest and we thus also include the case ${\omega}(D)$ being equal to $\frac{2}{3}(\Delta+1)$.
@@
Let $D$ be a connected digraph and $\Delta$ be an integer with $\Delta_{\max}(D)\leq\Delta$ and ${\omega}(D)\geq\frac{2}{3}(\Delta+1)$. Then $D$ has an acyclic set of vertices $I$ such that ${\omega}(D-I)=\overset{\text{\tiny$\bm{\leftrightarrow}$}}{\omega}(D)-1$, unless $D=\overleftrightarrow{C_{n}\circ K_{p}}$ with $n\geq 5$ odd and $p=\frac{1}{2}\overset{\text{\tiny$\bm{\leftrightarrow}$}}{\omega}(D)=\frac{1}{3}(\Delta+1)$.
We first prove the following result, which generalises to digraphs a more precise version of Haxell’s result obtained by King [23, Theorem 5].
@@
Let $k$ be a positive integer and $D$ be a digraph with vertices partitioned into bicliques $C_{1},\dots,C_{r}$. If for every $i\in[r]$ and every $v\in C_{i}$, $v$ has at most $k$ out-neighbours outside $C_{i}$ and at most $|C_{i}|-k$ in-neighbours outside $C_{i}$, then $D$ contains an acyclic set of size $r$.
Given a digraph $D$ whose vertices are partitionned into independent sets $V_{1},\dots,V_{r}$, an acyclic system of representative (with respect to $V_{1},\dots,V_{r}$), or ASR for conciseness, is an acyclic set of vertices $X$ such that, for every $i\in[r]$, $|V_{i}\cap X|=1$. If we relax this condition to $|V_{i}\cap X|\leq 1$, then $X$ is a partial ASR. Theorem 4.3 can be restated in terms of ASR as follows.
@@
Let $k$ be a positive integer and $D$ be a digraph with vertices partitioned into independent sets $V_{1},\dots,V_{r}$. If every vertex $v\in V_{i}$ satisfies $d^{+}(v)\leq k$ and $d^{-}(v)\leq|V_{i}|-k$, then $D$ admits an ASR.
@@
We proceed by induction on $r\geq 1$. The result being trivial when $r=1$, we assume $r\geq 2$. Let us fix any vertex $x_{1}\in V_{r}$, we will show as a stronger result that there exists an ASR containing $x_{1}$. Assume for a contradiction that no such ASR exists.
@@
    every vertex in $Y$ has exactly one in-neighbour in $X$ and every vertex in $X$ has at least one out-neighbour in $Y$, in particular $|X|\leq|Y|$ ; and
@@
Proof of claim. Assume for a contradiction that $(I,X,Y)$ is a good triplet of $D$. The assumption on the degrees in $D$ implies in particular $d^{+}(x)\leq k$ for every vertex $x\in X$ and $d^{-}(y)\leq|V_{\iota(y)}|-k$ for every vertex $y\in Y$. Observe that $|X|\leq|Y|=|I|$ by 3 and 2. We thus have
$\sum_{x\in X}d^{+}(x)+\sum_{y\in Y}d^{-}(y)\leq k|I|+\sum_{i\in I}(|V_{i}|-k)=|V_{I}|.$
@@
Our goal is now to show that $D$ admits an infinite sequence $(Y_{i})_{i\geq 1}$ of partial ASR, such that each $Y_{i}$ is a proper subset of $Y_{i+1}$. This will yield the contradiction since such a sequence has length at most $r$.
Let $R_{1}$ be an ASR of $D[V_{[r-1]}]$ (with respect to $V_{1},\dots,V_{r-1}$) for which the size of $Y_{1}=Y_{1}^{\prime}=N^{+}(x_{1})\cap R_{1}$ is minimum. The existence of $R_{1}$ is guaranteed by induction on $r$. Observe that $Y_{1}\neq\varnothing$ for otherwise $R_{1}\cup\{x_{1}\}$ is an ASR of $D$ containing $x_{1}$. We let $X_{1}$ be the singleton $\{x_{1}\}$. We will recursively build $(R_{i+1},X_{i+1},Y_{i+1},Y_{i+1}^{\prime})$ from $(R_{i},X_{i},Y_{i},Y_{i}^{\prime})$ in such a way that, for every $i\geq 1$, each of the following properties holds.
  1. 1.
@@
  2. 2.
    $R_{i}$ is an ASR of $D[V_{[r-1]}]$ such that for every $11\leq j\leq i-$ 1, $R_{i}\cap N^{+}(x_{j})=Y_{j}^{\prime}$. Subject to this property, $Y_{i}^{\prime}=N^{+}(x_{i})\cap R_{i}$ has minimum size.
  3. 3.
    $Y_{i}=\bigsqcup_{1\leq j\leq i}Y_{j}^{\prime}$ and $Y_{i}^{\prime}\neq\varnothing$.
Observe that $(R_{1},X_{1},Y_{1},Y_{1}^{\prime})$ satisfies 12, and 3. Also, the existence of such an infinite sequence $(R_{i},X_{i},Y_{i},Y_{i}^{\prime})_{i\geq 1}$ implies the desired infinite sequence $(Y_{i})_{i\geq 1}$ of partial ASR, yielding the contradiction.
Hence we assume that for some $i\geq 1$, there exists $(R_{i},X_{i},Y_{i},Y_{i}^{\prime})_{1\leq j\leq i}$ such that, for each $j\in[i]$12, and 3 hold. We construct $(R_{i+1},X_{i+1},Y_{i+1},Y_{i+1}^{\prime})$ as follows.
@@
[…]
    If such a vertex $x_{i+1}$ exists then $X_{i+1}=\{x_{1},\dots,x_{i+1}\}$ is acyclic as $x_{i+1}$ is a source in $D[X_{i+1}]$ and $X_{i}=\{x_{1},\dots,x_{i}\}$ is acyclic by recurrence, so 1 holds.
    Let us justify the existence of $x_{i+1}$. If no such vertex exists, then $X_{i}$ and $Y_{i}$ are two disjoint acyclic sets such that $x_{1}\in X_{i}$, with $Y_{i}$ being an ASR of $V_{\iota(Y_{i})}$. Combining conditions 2 and 3 we obtain that $(N^{+}(x_{1}),\dots,N^{+}(x_{i}))$ partitions $Y_{i}$ into non-empty parts, so in particular every vertex in $Y_{i}$ has exactly one in-neighbour in $X_{i}$ and every vertex in $X_{i}$ has at least one out-neighbour in $Y_{i}$. Hence, if no vertex in $V_{\iota(Y_{i})}\setminus(X_{i}\cup Y_{i})$ satisfies ($\star$4), we deduce that $(\iota(Y_{i}),X_{i},Y_{i})$ is a good triplet, a contradiction to Claim 4.6.
  • We let $R_{i+1}$ be an ASR of $D[V_{[r-1]}]$ such that for every $1\leq j\leq i$, $R_{i+1}\cap N^{+}(x_{j})=Y_{j}^{\prime}$, and subject to this property, $Y_{i+1}^{\prime}=N^{+}(x_{i+1})\cap R_{i+1}$ has minimum size. The existence of $R_{i+1}$ is guaranteed as $R_{i}$ is such an ASR. Hence 2 is maintained.
  • It remains to show that 3 is maintained, that is $Y_{i+1}^{\prime}\neq\varnothing$ and $Y_{i+1}^{\prime}\cap Y_{i}=\varnothing$. First we have $Y_{i+1}^{\prime}\cap Y_{i}=\varnothing$ by ($\star$4). Now assume for a contradiction that $Y_{i+1}^{\prime}=\varnothing$, which together with ($\star$4) implies that $x_{i+1}$ has no out-neighbour in $R_{i+1}$.
    By recurrence we know that $(Y_{1}^{\prime},\dots,Y_{i}^{\prime})$ partitions $Y_{i}$. As $x_{i+1}$ belongs to $V_{\iota(Y_{i})}\setminus(X_{i}\cup Y_{i})$ there exists a unique $j\leq i$ such that $\iota(x_{i+1})\in\iota(Y_{j}^{\prime})$. As $Y_{j}^{\prime}$ is a partial ASR, there exists a unique $y_{j}\in Y_{j}^{\prime}\cap V_{\iota(x_{i+1})}$. Consider the set $\tilde{R}_{j}$ defined as
@@
    Since $x_{i+1}$ has no in-neighbour in $X_{i}$ by ($\star$4), for every $1\leq\ell<j$, we have $\tilde{R}_{j}\cap N^{+}(x_{\ell})=R_{i+1}\cap N^{+}(x_{\ell})=Y_{\ell}^{\prime}$. We finally have $\tilde{R}_{j}\cap N^{+}(x_{j})=(R_{i+1}\cap N^{+}(x_{j}))\setminus\{y_{j}\}=Y_{j}^{\prime}\setminus\{y_{j}\}$, a contradiction to the choice of $R_{j}$.
    Hence 3 is maintained and the proof follows.∎
With Theorem 4.3 in hand, we are now going to prove Theorem 4.2. We need two additional intermediate results. Let $G$ be a (di)graph. We denote by $\mathcal{C}(G)$ the set of all maximum (bi)cliques of $G$. Given a collection $\mathcal{C}\subseteq\mathcal{C}(G)$ of maximum (bi)cliques of $G$, we denote by $\operatorname{IG}(\mathcal{C})$ the intersection graph of $\mathcal{C}$, i.e. the undirected graph with vertex-set $\mathcal{C}$ which contains an edge between $C_{1}$ and $C_{2}$ whenever $C_{1}\cap C_{2}\neq\varnothing$. We first need the following result due to Hajnal [17].
@@
$\left|\bigcap\mathcal{C}\right|+\left|\bigcup\mathcal{C}\right|\geq 2\omega(G).$
The second intermediate result we need is an extension of a result of Kostochka [25] due to Christofides, Edwards, and King [10]. The statement we use slightly differs from the original one (see [32, Lemma 6.12] for a proof of the result in the following form).
@@
Let $G$ be a graph with $\omega(G)\geq\frac{2}{3}(\Delta(G)+1)$, let $\mathcal{C}\subseteq\mathcal{C}(G)$ such that $\operatorname{IG}(\mathcal{C})$ is connected, let $X=\bigcap\mathcal{C}$ and let $Y=\bigcup\mathcal{C}$. If $X=\varnothing$ then $\omega(G)=\frac{2}{3}(\Delta(G)+1)$ and $Y$ has a partition $\{Q_{i}\mid i\in[n]\}$ into $n\geq 4$ cliques of $G$ each of cardinality $\frac{1}{2}\omega(G)$ such that $\mathcal{C}\setminus\{Q_{1}\cup Q_{n}\}=\{Q_{i}\cup Q_{i+1}\mid i\in[n-1]\}$.
@@
Let $D$ be a minimum counterexample, which means that for some integer $\Delta$, $\Delta_{\max}(D)\leq\Delta$, ${\omega}(D)\geq\frac{2}{3}(\Delta+1)$, $D$ is distinct from the obstruction $\overleftrightarrow{C_{n}\circ K_{p}}$ and $D$ does not contain any acyclic set of vertices $I$ intersecting every maximum biclique of $D$. We first show that every vertex $v$ belongs to a maximum biclique of $D$. Assume that this is not the case for some vertex $v$ and let $D^{\prime}$ be $D-v$. Observe that ${\omega}(D^{\prime})=\overset{\text{\tiny$\bm{\leftrightarrow}$}}{\omega}(D)\geq\frac{2}{3}(\Delta+1)\geq\Delta_{\max}(D^{\prime})$. By minimality of $D$, either $D^{\prime}$ has an acyclic set $I$ such that ${\omega}(D^{\prime}-I)=\overset{\text{\tiny$\bm{\leftrightarrow}$}}{\omega}(D^{\prime})-1$ or $D^{\prime}$ has a connected component $H=\overleftrightarrow{C_{n}\circ K_{p}}$ with $n\geq 5$ and $p=\frac{1}{2}\overset{\text{\tiny$\bm{\leftrightarrow}$}}{\omega}(D)$. In the former case, as $v$ does not belong to a maximum clique of $D$, $I$ is also an acyclic set of $D$ such that ${\omega}(D-I)=\overset{\text{\tiny$\bm{\leftrightarrow}$}}{\omega}(D)-1$, a contradiction to $D$ being a counterexample. In the latter case, every vertex $x\in V(H)$ satisfies $d^{+}_{H}(x)=d^{-}_{H}(x)=\Delta$. By connectivity of $D$, some vertex $x\in V(H)$ is adjacent to $v$, implying that $\max(d^{+}_{D}(x),d^{-}_{D}(x))>\Delta$, a contradiction to $\Delta_{\max}(D)\leq\Delta$.
@@
Observe that $v_{i}$ has at least ${\omega}(D)-1$ out-neighbours in $Y_{i}=\bigcup_{C\in V(\mathcal{H}_{i})}C$, so the out-degree of $v_{i}$ outside $X_{i}$ in $D^{\prime}$ is at most $\Delta-\overset{\text{\tiny$\bm{\leftrightarrow}$}}{\omega}(D)+1\leq\frac{1}{3}(\Delta+1)$. Since it is an integer, the out-degree of $v_{i}$ outside $X_{i}$ in $D^{\prime}$ is thus at most $k$. On the other hand, $v_{i}$ has at most $\Delta-|Y_{i}|+1$ in-neighbours outside $X_{i}$ in $D^{\prime}$. By Lemma 4.7 applied to $S(D)$ we have $|Y_{i}|\geq 2\overset{\text{\tiny$\bm{\leftrightarrow}$}}{\omega}(D)-|X_{i}|\geq\frac{4}{3}(\Delta+1)-|X_{i}|$. Hence $v_{i}$ has at most $|X_{i}|-\frac{1}{3}(\Delta+1)\leq|X_{i}|-k$ in-neighbours outside $X_{i}$ in $D^{\prime}$.
@@
Therefore, we can apply Theorem 4.3 to $D^{\prime}$, and conclude that $D^{\prime}$ admits an acyclic set $I$ such that ${\omega}(D^{\prime})=\overset{\text{\tiny$\bm{\leftrightarrow}$}}{\omega}(D-I)$. By construction of $D^{\prime}$, $I$ is also an acyclic set of $D$ such that ${\omega}(D)=\overset{\text{\tiny$\bm{\leftrightarrow}$}}{\omega}(D-I)$, a contradiction. $\lozenge$

### v3

@@ 5 Hitting every maximum biclique with an acyclic set of vertices
Let $G$ be a graph with maximum degree $\Delta(G)\leqslant 3k-1$. Haxell [22] proved that, if the vertices of $G$ can be partitioned into $r$ cliques $C_{1},\dots,C_{r}$ of size $\geqslant 2k$, then $G$ contains an independent set $S$ of size $r$. In particular, $S$ intersects each clique $C_{i}$.
Using this result, Rabern [33] proved that every graph $G$ satisfying $\omega(G)\geqslant\frac{3}{4}(\Delta(G)+1)$ contains an independent set intersecting every maximum clique of $G$. This was later improved to $\omega(G)>\frac{2}{3}(\Delta(G)+1)$ by King [25], which is best possible. Christofides, Edwards, and King [10] later characterised exactly the graphs $G$ satisfying $\omega(G)=\frac{2}{3}(\Delta(G)+1)$ that do not admit an independent set intersecting every maximum clique. Combined with King’s earlier result, they thus obtained the following. Given two graphs $G$ and $G^{\prime}$, we denote by $G\circ G^{\prime}$ the lexicographic product of $G$ by $G^{\prime}$, which is the graph with vertex-set $V(G)\times V(G^{\prime})$ and edge-set $\{(u,u^{\prime})(v,v^{\prime})\mid uv\in E(G)\textbf{~or~}u=v\text{~and~}u^{\prime}v^{\prime}\in E(G^{\prime})\}$.
Theorem 5.1 (Christofides et al. [10, Theorem 2]).
Let $G$ be a connected graph with $\omega(G)\geqslant\frac{2}{3}(\Delta(G)+1)$. Then $G$ has an independent set $I$ such that $\omega(G-I)=\omega(G)-1$, unless $G=C_{n}\circ K_{p}$ with $n\geqslant 5$ odd and $p=\frac{1}{2}\omega(G)$.
The goal of this section is to prove the following generalisation of Theorem 5.1. Note that, in the proof of Theorem 1.7, we will only use Theorem 5.2 with $\overset{\text{\tiny$\bm{\leftrightarrow}$}}{\omega}(D)$ being larger than $\frac{2}{3}(\Delta+1)$. We find the characterisation of the equality case of independent interest and we thus also include the case $\overset{\text{\tiny$\bm{\leftrightarrow}$}}{\omega}(D)$ being equal to $\frac{2}{3}(\Delta+1)$.
@@
Let $D$ be a connected digraph and $\Delta$ be an integer such that $\Delta_{\max}(D)\leqslant\Delta$ and $\overset{\text{\tiny$\bm{\leftrightarrow}$}}{\omega}(D)\geqslant\frac{2}{3}(\Delta+1)$. Then $D$ has an acyclic set of vertices $I$ such that $\overset{\text{\tiny$\bm{\leftrightarrow}$}}{\omega}(D-I)=\overset{\text{\tiny$\bm{\leftrightarrow}$}}{\omega}(D)-1$, unless $D=\overleftrightarrow{C_{n}\circ K_{p}}$ with $n\geqslant 5$ odd and $p=\frac{1}{2}\overset{\text{\tiny$\bm{\leftrightarrow}$}}{\omega}(D)=\frac{1}{3}(\Delta+1)$.
We first prove the following result, which generalises to digraphs a more precise version of Haxell’s result obtained by King [25, Theorem 5].
@@
Let $k$ be a positive integer and $D$ be a digraph with vertices partitioned into bicliques $C_{1},\dots,C_{r}$. If for every $i\in[r]$ and every $\in C_{i}, vv$ has at most $k$ out-neighbours outside $C_{i}$ and at most $|C_{i}|-k$ in-neighbours outside $C_{i}$, then $D$ contains an acyclic set of size $r$.
Given a digraph $D$ whose vertices are partitioned into independent sets $V_{1},\dots,V_{r}$, an acyclic system of representatives (with respect to $V_{1},\dots,V_{r}$), or ASR for conciseness, is an acyclic set of vertices $X$ such that, for every $i\in[r]$, $|V_{i}\cap X|=1$. If we relax this condition to $|V_{i}\cap X|\leqslant 1$, then $X$ is a partial ASR.
Note that, in the statement of Theorem 5.3, any acyclic set $X$ of size $r$ is necessarily a transversal of $(C_{1},\dots,C_{r})$, as two vertices of $X$ cannot belong to the same biclique. Therefore, Theorem 5.3 can be restated in terms of ASR as follows.
@@
Let $k$ be a positive integer and $D$ be a digraph with vertices partitioned into independent sets $V_{1},\dots,V_{r}$. If every vertex $v\in V_{i}$ satisfies $d^{+}(v)\leqslant k$ and $d^{-}(v)\leqslant|V_{i}|-k$, then $D$ admits an ASR.
@@
We proceed by induction on $r\geqslant 1$. The result being trivial when $r=1$, we assume $r\geqslant 2$. Let us fix any vertex $x_{1}\in V_{r}$, we will show as a stronger result that there exists an ASR containing $x_{1}$. Assume for a contradiction that no such ASR exists.
@@
    every vertex in $Y$ has exactly one in-neighbour in $X$ and every vertex in $X$ has at least one out-neighbour in $Y$, in particular $|X|\leqslant|Y|$ ; and
@@
Proof of claim. Assume for a contradiction that $(I,X,Y)$ is a good triplet of $D$. The assumption on the degrees in $D$ implies in particular $d^{+}(x)\leqslant k$ for every vertex $x\in X$ and $d^{-}(y)\leqslant|V_{\iota(y)}|-k$ for every vertex $y\in Y$. Observe that $|X|\leqslant|Y|=|I|$ by 3 and 2. We thus have
$\sum_{x\in X}d^{+}(x)+\sum_{y\in Y}d^{-}(y)\leqslant k|I|+\sum_{i\in I}(|V_{i}|-k)=|V_{I}|.$
@@
Our goal is now to show that $D$ admits an infinite sequence $(Y_{i})_{i\geqslant 1}$ of partial ASR, such that each $Y_{i}$ is a proper subset of $Y_{i+1}$. This will yield the contradiction since such a sequence has length at most $r$.
Let $R_{1}$ be an ASR of $D[V_{[r-1]}]$ (with respect to $V_{1},\dots,V_{r-1}$) for which the size of $Y_{1}=Y_{1}^{\prime}=N^{+}(x_{1})\cap R_{1}$ is minimum. The existence of $R_{1}$ is guaranteed by induction on $r$. Observe that $Y_{1}\neq\varnothing$ for otherwise $R_{1}\cup\{x_{1}\}$ is an ASR of $D$ containing $x_{1}$. We let $X_{1}$ be the singleton $\{x_{1}\}$. We will recursively build $(R_{i+1},X_{i+1},Y_{i+1},Y_{i+1}^{\prime})$ from $(R_{i},X_{i},Y_{i},Y_{i}^{\prime})$ in such a way that, for every $i\geqslant 1$, each of the following properties holds.
  1. (1)
@@
  2. (2)
    $R_{i}$ is an ASR of $D[V_{[r-1]}]$ such that for every $1\leqslant j\leqslant i-1$, $R_{i}\cap N^{+}(x_{j})=Y_{j}^{\prime}$. Subject to this property, $Y_{i}^{\prime}=N^{+}(x_{i})\cap R_{i}$ has minimum size.
  3. (3)
    $Y_{i}=\bigsqcup_{1\leqslant j\leqslant i}Y_{j}^{\prime}$ and $Y_{i}^{\prime}\neq\varnothing$.
Observe that $(R_{1},X_{1},Y_{1},Y_{1}^{\prime})$ satisfies (1)(2), and (3). Also, the existence of such an infinite sequence $(R_{i},X_{i},Y_{i},Y_{i}^{\prime})_{i\geqslant 1}$ implies the desired infinite sequence $(Y_{i})_{i\geqslant 1}$ of partial ASR, yielding the contradiction.
Hence we assume that for some $i\geqslant 1$, there exists $(R_{i},X_{i},Y_{i},Y_{i}^{\prime})_{1\leqslant j\leqslant i}$ such that, for each $j\in[i]$(1)(2), and (3) hold. We construct $(R_{i+1},X_{i+1},Y_{i+1},Y_{i+1}^{\prime})$ as follows.
@@
    $N^{-}(x_{i+1})\cap X_{i}=\varnothing\hskip 22.76228pt\text{and}\hskip 22.76228ptN^{+}(x_{i+1})\cap Y_{i}=\varnothing.$ ($\star$)
    If such a vertex $x_{i+1}$ exists then $X_{i+1}=\{x_{1},\dots,x_{i+1}\}$ is acyclic as $x_{i+1}$ is a source in $D[X_{i+1}]$ and $X_{i}=\{x_{1},\dots,x_{i}\}$ is acyclic by recurrence, so (1) holds.
    Let us justify the existence of $x_{i+1}$. If no such vertex exists, then $X_{i}$ and $Y_{i}$ are two disjoint acyclic sets such that $x_{1}\in X_{i}$, with $Y_{i}$ being an ASR of $V_{\iota(Y_{i})}$. Combining conditions (2) and (3) we obtain that $(N^{+}(x_{1}),\dots,N^{+}(x_{i}))$ partitions $Y_{i}$ into non-empty parts, so in particular every vertex in $Y_{i}$ has exactly one in-neighbour in $X_{i}$ and every vertex in $X_{i}$ has at least one out-neighbour in $Y_{i}$. Hence, if no vertex in $V_{\iota(Y_{i})}\setminus(X_{i}\cup Y_{i})$ satisfies ($\star$5), we deduce that $(\iota(Y_{i}),X_{i},Y_{i})$ is a good triplet, a contradiction to Claim 5.6.
  • We let $R_{i+1}$ be an ASR of $D[V_{[r-1]}]$ such that for every $1\leqslant j\leqslant i$, $R_{i+1}\cap N^{+}(x_{j})=Y_{j}^{\prime}$, and subject to this property, $Y_{i+1}^{\prime}=N^{+}(x_{i+1})\cap R_{i+1}$ has minimum size. The existence of $R_{i+1}$ is guaranteed as $R_{i}$ is such an ASR. Hence (2) is maintained.
  • It remains to show that (3) is maintained, that is $Y_{i+1}^{\prime}\neq\varnothing$ and $Y_{i+1}^{\prime}\cap Y_{i}=\varnothing$. First we have $Y_{i+1}^{\prime}\cap Y_{i}=\varnothing$ by ($\star$5). Now assume for a contradiction that $Y_{i+1}^{\prime}=\varnothing$, which together with ($\star$5) implies that $x_{i+1}$ has no out-neighbour in $R_{i+1}$.
    By recurrence we know that $(Y_{1}^{\prime},\dots,Y_{i}^{\prime})$ partitions $Y_{i}$. As $x_{i+1}$ belongs to $V_{\iota(Y_{i})}\setminus(X_{i}\cup Y_{i})$ there exists a unique $j\leqslant i$ such that $\iota(x_{i+1})\in\iota(Y_{j}^{\prime})$. As $Y_{j}^{\prime}$ is a partial ASR, there exists a unique $y_{j}\in Y_{j}^{\prime}\cap V_{\iota(x_{i+1})}$. Consider the set $\tilde{R}_{j}$ defined as
@@
    Since $x_{i+1}$ has no in-neighbour in $X_{i}$ by ($\star$5), for every $1\leqslant\ell<j$, we have $\tilde{R}_{j}\cap N^{+}(x_{\ell})=R_{i+1}\cap N^{+}(x_{\ell})=Y_{\ell}^{\prime}$. We finally have $\tilde{R}_{j}\cap N^{+}(x_{j})=(R_{i+1}\cap N^{+}(x_{j}))\setminus\{y_{j}\}=Y_{j}^{\prime}\setminus\{y_{j}\}$, a contradiction to the choice of $R_{j}$.
    Hence (3) is maintained and the proof follows.∎
With Theorem 5.3 in hand, we are now going to prove Theorem 5.2. We need two additional intermediate results. Let $G$ be a (di)graph. We denote by $\mathcal{C}(G)$ the set of all maximum (bi)cliques of $G$. Given a collection $\mathcal{C}\subseteq\mathcal{C}(G)$ of maximum (bi)cliques of $G$, we denote by $\operatorname{IG}(\mathcal{C})$ the intersection graph of $\mathcal{C}$, i.e., the undirected graph with vertex-set $\mathcal{C}$ which contains an edge between $C_{1}$ and $C_{2}$ whenever $C_{1}\cap C_{2}\neq\varnothing$. We first need the following result due to Hajnal [18].
@@
$\left|\bigcap\mathcal{C}\right|+\left|\bigcup\mathcal{C}\right|\geqslant 2\omega(G).$
The second intermediate result we need is an extension of a result of Kostochka [27] due to Christofides, Edwards, and King [10]. The statement we use slightly differs from the original one (see [35, Lemma 6.12] for a proof of the result in the following form).
@@
Let $G$ be a graph with $\omega(G)\geqslant\frac{2}{3}(\Delta(G)+1)$, let $\mathcal{C}\subseteq\mathcal{C}(G)$ such that $\operatorname{IG}(\mathcal{C})$ is connected, let $X=\bigcap\mathcal{C}$ and let $Y=\bigcup\mathcal{C}$. If $X=\varnothing$ then $\omega(G)=\frac{2}{3}(\Delta(G)+1)$ and $Y$ has a partition $\{Q_{i}\mid i\in[n]\}$ into $n\geqslant 4$ cliques of $G$ each of cardinality $\frac{1}{2}\omega(G)$ such that $\mathcal{C}\setminus\{Q_{1}\cup Q_{n}\}=\{Q_{i}\cup Q_{i+1}\mid i\in[n-1]\}$.
@@
Let $D$ be a minimum counterexample, which means that for some integer $\Delta$, $\Delta_{\max}(D)\leqslant\Delta$, $\overset{\text{\tiny$\bm{\leftrightarrow}$}}{\omega}(D)\geqslant\frac{2}{3}(\Delta+1)$, $D$ is distinct from the obstruction $\overleftrightarrow{C_{n}\circ K_{p}}$ and $D$ does not contain any acyclic set of vertices $I$ intersecting every maximum biclique of $D$.
We first show that every vertex $v$ belongs to a maximum biclique of $D$. Assume that this is not the case for some vertex $v$ and let $D^{\prime}$ be DvD-v. Observe then that $\overset{\text{\tiny$\bm{\leftrightarrow}$}}{\omega}(D^{\prime})=\overset{\text{\tiny$\bm{\leftrightarrow}$}}{\omega}(D)\geqslant\frac{2}{3}(\Delta+1)\geqslant\Delta_{\max}(D^{\prime})$. By minimality of $D$, either $D^{\prime}$ has an acyclic set $I$ such that $\overset{\text{\tiny$\bm{\leftrightarrow}$}}{\omega}(D^{\prime}-I)=\overset{\text{\tiny$\bm{\leftrightarrow}$}}{\omega}(D^{\prime})-1$ or $D^{\prime}$ has a connected component $H=\overleftrightarrow{C_{n}\circ K_{p}}$ with $n\geqslant 5$ and $p=\frac{1}{2}\overset{\text{\tiny$\bm{\leftrightarrow}$}}{\omega}(D)$. In the former case, as $v$ does not belong to a maximum clique of $D$, II is also an acyclic set of $D$ such that $\overset{\text{\tiny$\bm{\leftrightarrow}$}}{\omega}(D-I)=\overset{\text{\tiny$\bm{\leftrightarrow}$}}{\omega}(D)-1$, a contradiction to $D$ being a counterexample. In the latter case, every vertex $x\in V(H)$ satisfies $d^{+}_{H}(x)=d^{-}_{H}(x)=\Delta$. By connectivity of $D$, some vertex $x\in V(H)$ is adjacent to $v$, implying that $\max(d^{+}_{D}(x),d^{-}_{D}(x))>\Delta$, a contradiction to $\Delta_{\max}(D)\leqslant\Delta$.
@@
Observe that $v_{i}$ has at least $\overset{\text{\tiny$\bm{\leftrightarrow}$}}{\omega}(D)-1$ out-neighbours in $Y_{i}=\bigcup_{C\in V(\mathcal{H}_{i})}C$, so the out-degree of $v_{i}$ outside $X_{i}$ in $D^{\prime}$ is at most $\Delta-\overset{\text{\tiny$\bm{\leftrightarrow}$}}{\omega}(D)+1\leqslant\frac{1}{3}(\Delta+1)$. Since it is an integer, the out-degree of $v_{i}$ outside $X_{i}$ in $D^{\prime}$ is thus at most $k$. On the other hand, $v_{i}$ has at most $\Delta-|Y_{i}|+1$ in-neighbours outside $X_{i}$ in $D^{\prime}$. By Lemma 5.7 applied to $S(D)$ we have $|Y_{i}|\geqslant 2\overset{\text{\tiny$\bm{\leftrightarrow}$}}{\omega}(D)-|X_{i}|\geqslant\frac{4}{3}(\Delta+1)-|X_{i}|$. Hence $v_{i}$ has at most $|X_{i}|-\frac{1}{3}(\Delta+1)\leqslant|X_{i}|-k$ in-neighbours outside $X_{i}$ in $D^{\prime}$.
@@
Therefore, we can apply Theorem 5.3 to $D^{\prime}$, and conclude that $D^{\prime}$ admits an acyclic set $I$ of size $r$, which is thus a transversal of $(X_{1},\dots,X_{r})$. By definition of $D^{\prime}$ and $(X_{1},\dots,X_{r})$, $I$ is thus an acyclic set of $D$ such that $\overset{\text{\tiny$\bm{\leftrightarrow}$}}{\omega}(D-I)=\overset{\text{\tiny$\bm{\leftrightarrow}$}}{\omega}(D)-1$, a contradiction. $\lozenge$